you are given three stock cultures- E coli at a concentration of 1.5 x 10^3 cfu/ml, Bacillus at a concentration of 4.2 x10^3 cfu/ml, and staphylococcus at a concentratino of 2.4 x10^4 cful/ml. you ahve been asked to make a 30 ml of a single mixed culture with a final concentration of 500 cfu/ml of E coli, 150 cfu/ml of bacillus, and 80 cfu/ml of Staph. How would you make this mixture?
Answer #1

E coli concentration of 1.5 x 10^3 cfu/ml

Bacillus concentration of 4.2 x10^3 cfu/ml

staphylococcus concentratino of 2.4 x10^4 cful/ml

to make 30ml of 500 cfu/ml of E coli, 150 cfu/ml of bacillus, and 80 cfu/ml of Staph following steps have to be taken

It will follow simple molarity rule

M1*V1=M2*V2

for E.coli, 1.5 x 10^3 * V1 = 500 * 30

V1 = 500*30/1.5X103

=10ml

therefore from the stock 10 ml of Ecoli will be taken

for bacillus to get 150 cfu/ml

M3 X V3= M4 X V4

4.2 x10^3 X V3 = 150 X 30

V3=150 X 30/4.2 x10^3

=1.07ml

therefore 1.07 ml of stock bacilli has to betaken from stock of 4.2 x10^3 cfu/ml.

For staph,

M5XV5=M6Xv6

2.4 x10^4 X V5= 80 X 30

V5 = 0.1ml

therefore to make 30 ml of the mixed culture, 10 ml of Ecoli, 1.07ml of Bacillus and 0.1ml of staph will be taken from the given stocks, and the rest of the volume that is 30 - 11.17= 18.83ml will be make up with autoclaved distilled water.